OpenStudy (anonymous):

The supplement of an angle is 36 degrees less than twice the supplement of the complement of the angle. Find the measure of the supplement.

OpenStudy (amistre64):

sounds political

OpenStudy (anonymous):

how is it a -?

OpenStudy (amistre64):

Sa = -36 + 2d d = c + (s-c) Sa = -36 + 2(c+(s-c))

OpenStudy (amistre64):

what?

OpenStudy (anonymous):

how is 36 a negative?

OpenStudy (amistre64):

it says so; 36 less than ....

OpenStudy (anonymous):

okay so is the sa thing is the answer?

OpenStudy (amistre64):

its what im calling the supplement to the angle; or Sa

OpenStudy (amistre64):

180 - 90 - a = 2(180 - (180-(90-a))-a) - 36, is what i think it reads as

OpenStudy (amistre64):

i think ive almost got it right :)

OpenStudy (anonymous):

okay well thanks for your help

OpenStudy (amistre64):

i cant seem to concentrate on it enough to determine whats what at the moment ...

OpenStudy (amistre64):

a = 12 .... pretty sure

OpenStudy (amistre64):